OpenStudy (anonymous):

z=5t+3i+t(t-4)i Determinate t so z=real numbers

OpenStudy (anonymous):

For z to be a real number, t(t-4)=-3

OpenStudy (anonymous):

Now, can u solve for t

OpenStudy (anonymous):

z= 3 and 1?

OpenStudy (anonymous):

Nope, solve for t

OpenStudy (anonymous):

so t=3 and 1

OpenStudy (anonymous):

t(t-4)=-3 t^2-4t+3=0

OpenStudy (anonymous):

YEP, t= 3 and 1

OpenStudy (anonymous):

Thanks ! =)

OpenStudy (anonymous):

:)yw